A rectangle has vertices Q(5,6), R(5,-3), S(1,-3), and T(1,6). Find
a) the perimeter
b) the area

Answer by checkley71(8403)

DISTANCE BETWEEN Q & R=6+3=9
DISTANCE BETWEEN S & T=6+3=9
DISTANCE BETWEEN Q & S=5-1=4
DISTANCE BETWEEN R & T=5-1=4
PERIMETER=2*9+2*4=18+8=26 ANSWER.
AREA=9*4=36 ANSWER.